c) A box contains 25 Balls, 20% of them are white whereas the rest of them are black. A person randomly 

chooses 3 balls from the box without replacement, find the probability of getting 3 black balls. 

 (6 marks) 

d) 20% of all Malaysian citizens are under the age of 15. A company randomly chooses 3 Malaysian 

citizens (without replacement), find the probability that all the three people chosen are age 15 or older.


1
Expert's answer
2020-08-03T18:23:21-0400

Solution c)


total balls = 25


white balls = 20% of 25 = 5

black balls = 25 - 5 = 20


Let probability of picking a black ball = p(B)


1st pick:


"p(B) = black balls \/ Total balls\n= 20 \/ 25 = 0.80"

Balls left:


black balls = 19

total balls = 24


2nd pick:


"p(B) = 19 \/ 24 \n= 0.79167"

Balls left:


black balls = 18

total balls = 23


3rd pick:


"p(B) = 18 \/ 23 \n= 0.78261"

Hence probability of picking 3 black balls without replacement:


"= 0.8 * 0.79167 * 0.78261"

Answer: 0.4957


Solution d)


population under 15 years = 20%

population 15 years or more = 80%


p(age > 15) = 0.8


This is the population proportion. Hence the probability of selecting a person without replacement remains the same for each trial


1st pick: p(age > 15) = 0.8

2nd pick: p(age > 15) = 0.8

3rd pick: p(age > 15) = 0.8


The probability that 3 people chosen from the population are 15 years or more:



"=0.8 * 0.8 * 0.8"

Answer: 0.512